OpenStudy (korosh23):

Pre-calculus 12 Exponential Equations Not Requiring Logarithms Solve the equation for x. 10^-3x . 10^x = 1/10

OpenStudy (korosh23):

\[10^{-3x}. 10^{x}= \frac{ 1 }{ 10}\]

hartnn (hartnn):

\(A^B\times A^C = A^{B+C}\) so how would you simplify the left side?

OpenStudy (korosh23):

Yes, same base, add exponents. 10^-2x= 1/10

OpenStudy (korosh23):

Do we consider 1/10 as 10^-1

hartnn (hartnn):

\(\dfrac{1}{x^n} = x^{-n}\) YES

OpenStudy (korosh23):

ok in that case x = -1/-2 = 1/2 The answer is correct. Thank you

hartnn (hartnn):

welcome ^_^